OpenStudy (anonymous):

you can afford monthly deposits of $210 into an account that pays 3.9% compounded monthly. how long will it be until you have 5800 to buy a boat?

OpenStudy (amistre64):

5800 = 210(1-k^(n*12))/(1-k), k=1+.039/12

OpenStudy (amistre64):

if we just let n be months, then there is no need for the n*12 bit

OpenStudy (amistre64):

B = P(1-k^n)/(1-k) B(1-k)/P = 1-k^n k^n = 1- B(1-k)/P n ln(k) = ln(1- B(1-k)/P) n = ln(1- B(1-k)/P)/ln(k)

OpenStudy (amistre64):

about 26 and a 1/2 months i believe

OpenStudy (cwrw238):

you are using sum of a geometric series amistre, right?

OpenStudy (amistre64):

yes, and trying to determine if this would amount to an annuity due; i believe ive figured it for an ordinary annuity

OpenStudy (anonymous):

thank you so much are you positive

OpenStudy (amistre64):

im iffy, but fairly sure :)

OpenStudy (amistre64):

m0 = 210 m1 = 210k + 210 m2 = 210k^2 + 210k + 210 mn = 210(geo sum)

OpenStudy (anonymous):

do you get the same answer if youuse this calculator i found online http://www.thecalculatorsite.com/finance/calculators/compoundinterestcalculator.php

OpenStudy (anonymous):

i typed things in and the calc isnt working idk

OpenStudy (amistre64):

S = 1 + k + k^2 + k^n -kS = -k -k^2 -k^n -k^(n+1) ------------------------------ S = (1 - k^(n+1))/(1-k)

OpenStudy (amistre64):

i believe i am 1 month heavy

OpenStudy (amistre64):

25 and a half months is what im wanting to go with ...

OpenStudy (anonymous):

thats what the calc gave u?

OpenStudy (amistre64):

no, thats what my own progression gave me

OpenStudy (anonymous):

okay sweet can you just chek that link i sent to see if it matches because when i put it in it says that imj missing something idk if itll work for u to make sure that 25 and a half months is the right answer

OpenStudy (amistre64):

youd have to put in a time period; say 4 years

OpenStudy (amistre64):

well, 30 months to be certain, start it at 0

OpenStudy (amistre64):

so youll see between 26 and 27 months the balance of 5800 slides in there

OpenStudy (anonymous):

ohhh so idk if its 26 or 27

OpenStudy (anonymous):

this is what i get for month 26 and 27 26 $ 210.00 $ 38.93 $ 11,260.00 $ 756.71 $ 12,016.71 27 $ 210.00 $ 39.74 $ 11,470.00 $ 796.45 $ 12,266.45

OpenStudy (amistre64):

well, 26 is too short, so if you have to do months; bump it up to 27

OpenStudy (anonymous):

okay thank u soooooooooo much

OpenStudy (anonymous):

could you pleas ehelp me with another question please i am going to post it now

OpenStudy (amistre64):

youre welcome; so i guess my 26 and a half would have been fine :)

OpenStudy (amistre64):

ive got to move along at the moment

OpenStudy (anonymous):

okay